# Predicting Critical Heat Flux with Uncertainty Quantification and Domain Generalization Using Conditional Variational Autoencoders and Deep Neural Networks

### v1

@@ Abstract
Deep generative models (DGMs) have proven to be powerful in generating realistic data samples. Their capability to learn the underlying distribution of a dataset enable them to generate synthetic data samples that closely resemble the original training dataset, thus addressing the challenge of data scarcity. In this work, we investigated the capabilities of DGMs by developing a conditional variational autoencoder (CVAE) model to augment the critical heat flux (CHF) measurement data that was used to generate the 2006 Groeneveld lookup table. To determine how this approach compared to traditional methods, a fine-tuned deep neural network (DNN) regression model was created and evaluated with the same dataset. Both the CVAE and DNN models achieved small mean absolute relative errors, with the CVAE model maintaining more favorable results. To quantify the uncertainty in the model’s predictions, uncertainty quantification (UQ) was performed with repeated sampling of the CVAE model and ensembling of the DNN model. Following UQ, the DNN ensemble notably improved performance when compared to the baseline DNN model, while the CVAE model achieved similar results to its non-UQ results. The CVAE model was shown to have significantly less variability and a higher confidence after assessment of the prediction-wise relative standard deviations. Evaluating domain generalization, both models achieved small mean error values when predicting both inside and outside the training domain, with predictions outside the training domain showing slightly larger errors. Overall, the CVAE model was comparable to the DNN regression model in predicting CHF values but with better uncertainty behavior.
@@ 1 Introduction
In this study, we developed a CVAE-based DGM for data augmentation of the NRC CHF dataset. One may wonder that $\sim$ 25,000 CHF data points seem to be a good amount of training data for ML tasks, why is DGM necessary in this case? We argue that after considering the wide ranges of reactor operating conditions, geometries and reactor types, the CHF data points are rather scarce and unbalanced, making data augmentation necessary, especially for heavily parameterized predictive ML models such as neural networks. In order to assess the performance of the CVAE generative model, we directly compared it with a regular deep neural network (DNN) model. DNN models have proven to be powerful for regression tasks by directly learning the mapping between inputs and outputs in the training data. Both the CVAE and fine-tuned DNN models were trained and tested using the same datasets to generate CHF data under the same specified conditions, with the results subsequently compared. Uncertainty quantification (UQ) analysis was then applied via sampling of the CVAE model and ensembling of the DNN model. Domain generalization analysis was also performed for both models to assess their ability in extrapolating to new, unseen domains.
Several metrics were applied to assess the performance of the models. Comparison with true holdout values showed close alignment, indicating a low level of error in the generated and predicted CHF values. UQ results showed that the CVAE model has very small uncertainties in the generation of CHF values, whereas higher values were observed for the DNN model. Both the CVAE and DNN models showed similar behavior in their ability to generalize beyond their training domain, achieving small mean absolute relative errors. Both models have proven to be reliable in generating and predicting CHF values outside the training domain. The mean relative absolute errors were slightly smaller when predicting within the training domain, compared to predicting outside the training domain. The findings from this work show that the CVAE and DNN models were successful at generating and predicting CHF values with a high degree of accuracy, with the CVAE demonstrating a significantly more favorable uncertainty behavior.
@@ 2 The CHF Experimental Dataset
CHF, also known as departure from nucleate boiling, is a critical phenomenon in heat transfer. It occurs when a heated surface reaches a point where it can no longer efficiently transfer heat to the surrounding fluid. In nuclear reactors, exceeding the CHF limit could potentially lead to fuel rod failure, which is a significant safety concern especially for pressurized water reactors. Therefore, it is essential to limit the heat flux of the fuel rods to a value below the CHF threshold. The collection of CHF measurement data can be challenging due to the nature of the experiments. Recently, the US NRC has published the largest known CHF dataset publicly available [24]. It consists of nearly 25,000 CHF measurements in vertical uniformly-heated water-cooled cylindrical tubes, gathered over a span of 60 years from 59 different sources. In these experiments, CHF values were measured at various TH initial/boundary conditions: pressure ($P$), mass flux ($G$), inlet temperature ($T_{\text{in}}$), as well as geometrical parameters like test section diameter ($D$), and heated length ($L$). Additionally, the dataset contains calculated parameters derived from measurements and water properties, including outlet equilibrium quality ($X$), and inlet enthalpy ($\Delta h_{\text{in}}$). The distributions of the values of these parameters, along with their pair-wise correlations, are illustrated in Figure 1.
Figure 1: The distributions and correlations of the TH conditions and CHF values in the NRC CHF dataset.
@@ 3.1 DNNs
In a regular feed-forward architecture, the mathematical expression of a DNN can be briefly defined as below:
  • Input layer:
    $\mathcal{Y}^{0}(\mathbf{x})=\mathbf{x}\in\mathbb{R}^{d_{\text{in}}}$ (1)
  • Hidden layer:
    $\mathcal{Y}^{l}(\mathbf{x})=\sigma\left(\mathbf{W}_{l}\mathcal{Y}^{l-1}(\mathbf{x})+\mathbf{b}_{l}\right)\in\mathbb{R}^{d_{l}},\text{for }1\leq l\leq L-1$ (2)
  • Output layer:
    $\mathcal{Y}^{L}(\mathbf{x})=\sigma\left(\mathbf{W}_{L}\mathcal{Y}^{L-1}(\mathbf{x})+\mathbf{b}_{L}\right)\in\mathbb{R}^{d_{\text{out}}}$ (3)
where $\mathbf{x}$ is the input vector with dimension $d_{\text{in}}$, $\mathcal{Y}^{l}(\mathbf{x})$ is the activation of the $l^{\text{th}}$ hidden layer with dimension $d_{l}$, and $\mathcal{Y}^{L}(\mathbf{x})$ is the prediction in the output layer with dimension $d_{\text{out}}$. $L$ is the depth of the neural network, including ($L-1$) hidden layers and one output layer. $\sigma$ is the activation function. $\mathbf{W}_{l}$ is the matrix of weights for the $l^{\text{th}}$ hidden layer. $\mathbf{b}_{l}$ is the vector of bias for the $l^{\text{th}}$ hidden layer.
@@ 4.1 The CVAE Generative Model
The CVAE generative model will be used to approximate the underlying distribution of the NRC CHF dataset to generate new CHF values under specific TH conditions. The vector of TH conditions consists of seven parameters: $P$, $G$, $T_{\text{in}}$, $D$, $L$, $X$, and $\Delta h_{\text{in}}$, as discussed in Section 2. During the training process, these seven parameters were provided to the model as conditions, that is, the vector $\mathbf{c}$ in Equation (5) and Figure 2, allowing it to learn how to generate CHF values under user desired conditions. If a conventional VAE generative model were to be trained, one would not condition the model on the vector of TH parameters. In this case, the trained VAE model can only generate synthetic samples randomly, making it difficult to evaluate its performance by comparing with the real experimental data, since it is almost impossible for the random samples to have TH conditions that exactly match the real experimental data.
The model architecture was optimized by performing hyperparameter tuning for 400 iterations. Both the encoder and decoder networks were constructed with four fully connected layers. Training was conducted over around 230 epochs with a batch size of 76. Prior to training, the data underwent shuffling and standardization. The model was trained on 80% of the NRC CHF dataset, while the remaining 20% was reserved for the testing and validation, where the validation set was used for hyperparameter tuning. TensorFlow [31] was used to build the CVAE model. Once the model was trained, the decoder was used to generate CHF values under specific TH conditions in the testing dataset. Subsequently, the generated CHF values were compared with the corresponding experimental CHF values in the testing dataset to evaluate their accuracy.
@@ 4.2 The DNN Predictive Model
The DNN predictive model was also constructed in TensorFlow using eight hidden layers to transform the $7$ input parameters listed in Section 4.1. The network depth, number of neurons per layer, activation functions, and learning rate were optimized using RayTune [32] using $1000$ hyperparameter configurations in a random search. This tuning process was performed using the validation partition of the dataset. The training, validation, and testing datasets used between the CVAE and DNN were identical to ensure an accurate and fair comparison between the two methods. Once the hyperparameters were set, the model was trained for a total of $500$ epochs with an exponential learning decay rate of $0.96$. The fully trained network was then evaluated against the testing set using the set of metrics provided in Section 5. The DNN model is much more straightforward to train since it only builds a black-box surrogate model of the training dataset without considering the data distribution. Our goal is to use such a fine tuned DNN model to “benchmark” the performance of the CVAE generative model using a comprehensive collection of metrics presented in the following sections.
@@ 4.3 UQ of the ML Models
Quantifying uncertainties in the generated samples by the CVAE model was done by utilizing the inherent variability in the generation process. Recall that the latent space in the CVAE model consists of two vectors that represent the mean value and standard deviation of the encoded distribution, which is usually assumed to be Gaussian. We first obtained 200 samples from the latent space distribution. Each of the latent vector samples is then combined with an input TH condition vector (the conditioning vector $\mathbf{c}$). The combined vector will be fed into the decoder to produce a CHF value. A different CHF value will be generated each time the decoder takes the same TH condition but different latent vector. This process was repeated for all TH conditions in the testing dataset. Afterwards, the mean $\mu_{\text{samples}}$ and standard deviation $\sigma_{\text{samples}}$ were computed among the 200 samples for each TH condition. $\sigma_{\text{samples}}$ will be used to represent the uncertainty in the generated samples. It also indicates how sensitive the model predictions are to changes in the latent vector.
@@ 4.4 Domain Generalization
An important factor in ML model evaluation is assessing the model’s ability to generalize to data outside the training domain. This involved assessing how accurately can the model extrapolate to new regions with unseen data. To this end, we employed the concept of a convex hull to evaluate both the CVAE and DNN models. The convex hull is the smallest convex set that includes all points in a given dataset. In order to determine whether the point we are predicting at is inside or outside the training domain, we first construct a convex hull based on the training dataset. Afterwards, the points in the testing dataset are classified into two categories; inside the convex hull or outside the convex hull. This is done by examining each testing data point to ascertain whether it falls within or outside the convex hull. If inside, the points are deemed to be within the domain covered by the training data; if outside, then the point is not represented in the space covered by the training data.
In this study, our testing dataset comprises 2,458 data points (10% of the NRC CHF dataset). Of these, 1,047 points fall within the convex hull defined by the training dataset, while 1,411 points lie outside. Evaluating the two models’ performance on generated/predicted CHF values for the 1,411 points outside the convex hull allows us to assess its capability for domain generalization.
@@ 5.1 Results of CHF Generation and Prediction using CVAE and DNN
Table 1 includes statistical metrics calculated based on the absolute relative errors and the $R^{2}$ metric for both models. It provides the mean, maximum, and standard deviations of the absolute relative error values, along with the fraction of testing points resulting in an absolute relative error value greater than 10%. The statistical metrics show that both models were able to predict CHF values accurately, with mean absolute relative error values of 1.4907% for the CVAE model and 1.8473% for the DNN model. The maximum error values were significantly higher than the mean values, however, it is reassuring to note that only 1.34% of the points in the testing data resulted in an absolute relative error greater than 10% for the DNN model compared to 0.56% for the CVAE model.
@@
$\mu_{\text{error}}$ 1.8473 % 1.4907 %
$\text{Max}_{\text{error}}$ 31.056 % 24.962 %
$\text{Std}_{\text{error}}$ 2.4065 % 1.8593 %
$F_{\text{error}}>10\%$ 1.3426 % 0.5695 %
$\text{R}^{2}$ 0.9990 0.9987
The small standard deviation values for both models indicate that the errors are not widely spread from the mean, suggesting that most predictions are close to the actual CHF values. Furthermore, both networks are observed with an $R^{2}$ value above 0.99 indicating a very high degree of correlation between true and generated/predicted values. Overall, these results show that both models perform well in generating and predicting CHF values, with the CVAE having more favorable values across all error metrics.
The last row in Figure 1 presents the pair-wise correlations between the measured CHF values and each of the seven parameters in the vector of TH conditions. It is desirable to maintain such correlations in the generated and predicted CHF data. The correlations between the predicted and generated CHF values and the TH parameters in the testing dataset were compared with the real data to determine whether the models can learn and preserve these correlations. Figure 4 compares the correlations of the TH parameters with the true CHF values and the generated CHF values using the CVAE model. Similarly, Figure 5 compares the correlations of the TH parameters with the true CHF values and the predicted CHF values using the DNN model. These figures also illustrate the TH parameter ranges in the testing dataset.
@@
Figures 4 and 5 show that the correlations between the TH parameters and the generated/predicted CHF values align closely with those found in the real data. Both models have consistent behavior in terms of the correlations. While largest relative absolute error is observed at smaller CHF values, the largest absolute deviations from the true values are observed at higher CHF values, which are associated with smaller values of $D$, $L$, $P$ and $T_{\text{in}}$, subcooled coolant conditions, and high $G$ values. It is important to note that these regions contain fewer data points in both training and testing datasets, resulting in the model’s deviations from the true values under these conditions. However, these deviations are not significant and do not heavily influence the error distributions, as these CHF conditions are less common in practical applications and thus in the CHF dataset.
@@ 5.2 Results of UQ
To quantify the model uncertainty, 200 unique samples were computed for each of the testing set’s input entries using the trained CVAE model. From these, the means and standard deviations of these samples were taken for each of these input entries. The ensemble approach described in Section 4.3 was then implemented using $n=20$ models each initialized with a different random seed. Each of these models were identical with the exception of the initialization. As with the CVAE model, the means and standard deviations were computed along the 200 samples for each input vector of the testing set. The standard deviations were then taken to compute the relative standard deviation using Equation 6. This metric standardizes the uncertainty measure by expressing the standard deviation as a percentage of the mean, making it easier to interpret and to compare across different scales of the outputs.
@@
This process resulted in a relative standard deviation value for each of the output means, providing quantification of the model’s uncertainty. By comparing the relative standard deviations from the CVAE to those of the ensemble of DNNs, we can assess the effectiveness of each approach in capturing the uncertainty associated with the model predictions. This is done using the 6 metrics reported in Table 2, which also include the standard deviation of the error across all prediction means, as was done in Section 5.1. In terms of error values, the DNN achieves a significantly smaller mean relative error of $0.8868\%$ when compared to the $1.4797\%$ of the CVAE. The standard deviation of this error distribution is also tighter in the DNN’s case, $1.5364\%$ versus $1.8496\%$, but with generally comparable values when considering their small magnitudes. The DNN’s maximum value of the error distribution, $37.307\%$, is observed to be significantly larger than the CVAE’s of $22.334\%$. Both of these maximum values are located at relatively small CHF values near the origin.
@@
$\mu_{\text{error}}$ 0.8868 % 1.4797 %
$\text{Max}_{\text{error}}$ 37.307 % 22.334 %
$\text{Std}_{\text{error}}$ 1.5364 % 1.8496%
Mean relative std 1.8023 % 0.2579 %
Max relative std 29.651 % 5.5714 %
$F_{\text{error}}>10\%$ 0.3662 % 0.5695 %
Now considering the distribution of relative standard deviations, the CVAE’s results report significantly smaller mean and maximum relative Std values compared to the DNN’s. The mean relative standard deviation of the DNN ($1.8023\%$) is over seven times greater than that of the CVAE ($0.2579\%$), indicating that the CVAE is more confident in its predictions with less variability in its outputs. This is similarly mirrored in their maximum values, where the DNN has a value of nearly $30\%$ with the CVAE reporting $5.5714\%$. The final metric, the fraction of predictions with relative errors above $10\%$, has comparable values between the two models. Considering the combination of the metrics in Table 2, the CVAE outputs are shown to have similar performance compared to a DNN ensemble while achieving model uncertainties significantly smaller.

### v2

@@ Abstract
Deep generative models (DGMs) can generate synthetic data samples that closely resemble the original dataset, addressing data scarcity. In this work, we developed a conditional variational autoencoder (CVAE) to augment critical heat flux (CHF) data used for the 2006 Groeneveld lookup table. To compare with traditional methods, a fine-tuned deep neural network (DNN) regression model was evaluated on the same dataset. Both models achieved small mean absolute relative errors, with the CVAE showing more favorable results. Uncertainty quantification (UQ) was performed using repeated CVAE sampling and DNN ensembling. The DNN ensemble improved performance over the baseline, while the CVAE maintained consistent results with less variability and higher confidence. Both models achieved small errors inside and outside the training domain, with slightly larger errors outside. Overall, the CVAE performed better than the DNN in predicting CHF and exhibited better uncertainty behavior.
@@ 1 Introduction
In this study, we develop a CVAE-based DGM for data augmentation of the NRC CHF dataset. One may wonder that $\sim$ 25,000 CHF data points seem to be a good amount of training data for ML tasks, why is DGM necessary in this case? We argue that after considering the wide ranges of reactor operating conditions, geometries and reactor types, the CHF data points are rather scarce and unbalanced, making data augmentation necessary, especially for heavily parameterized predictive ML models such as neural networks. In order to assess the performance of the CVAE generative model, we directly compared it with a regular deep neural network (DNN) model. DNN models have proven to be powerful for regression tasks by directly learning the mapping between inputs and outputs in the training data. Both the CVAE and fine-tuned DNN models were trained and tested using the same datasets to generate CHF data under the same specified conditions, with the results subsequently compared. Uncertainty quantification (UQ) analysis was then applied via sampling of the CVAE model and ensembling of the DNN model. Domain generalization analysis was also performed for both models to assess their ability in extrapolating to new, unseen domains.
Several metrics were applied to assess the performance of the models. Comparison with true holdout values showed close alignment, indicating a low level of error in the generated and predicted CHF values. UQ results showed that the CVAE model has very small uncertainties in the generation of CHF values, whereas higher values were observed for the DNN model. Both the CVAE and DNN models showed similar behavior in their ability to generalize beyond their training domain, achieving small mean absolute relative errors. Both models have proven to be reliable in generating and predicting CHF values outside the training domain. The mean relative absolute errors were slightly smaller when predicting within the training domain, compared to predicting outside the training domain. The findings from this work show that the CVAE and DNN models were successful at generating and predicting CHF values with a high degree of accuracy, with the CVAE demonstrating a more favorable error behavior.
@@ 2 The CHF Experimental Dataset
CHF, also known as departure from nucleate boiling, is a critical phenomenon in heat transfer. It occurs when a heated surface reaches a point where it can no longer efficiently transfer heat to the surrounding fluid. In nuclear reactors, exceeding the CHF limit could potentially lead to fuel rod failure, which is a significant safety concern especially for pressurized water reactors. Therefore, it is essential to limit the heat flux of the fuel rods to a value below the CHF threshold. The collection of CHF measurement data can be challenging due to the nature of the experiments. Recently, the US NRC has published the largest known CHF dataset publicly available [24]. It consists of nearly 25,000 CHF measurements in vertical uniformly-heated water-cooled cylindrical tubes, gathered over a span of 60 years from 59 different sources. In these experiments, CHF values were measured at various TH initial/boundary conditions: pressure ($P$), mass flux ($G$), inlet temperature ($T_{\text{in}}$), as well as geometrical parameters like test section diameter ($D$), and heated length ($L$). Additionally, the dataset contains calculated parameters derived from measurements and water properties, including outlet equilibrium quality ($x_{\text{e}}$), and inlet enthalpy ($\Delta h_{\text{in}}$). In this study, we selected $P$, $G$, $D$, $L$ and $x_{\text{e}}$ as the key TH parameters, along with CHF for the analysis. The distributions of the values of these parameters, along with their pair-wise correlations, are illustrated in Figure 1.
Figure 1: The distributions and correlations of the TH parameters and CHF values in the NRC CHF dataset.
@@ 4.1 The CVAE Generative Model
The CVAE generative model will be used to approximate the underlying distribution of the NRC CHF dataset to generate new CHF values under specific TH conditions. The vector of TH conditions consists of five parameters: $P$, $G$, $D$, $L$, and $x_{\text{e}}$, as discussed in Section 2. During the training process, these five parameters were provided to the model as conditions, that is, the vector $\mathbf{c}$ in Equation (2) and Figure 2, allowing it to learn how to generate CHF values under user desired conditions. If a conventional VAE generative model were to be trained, one would not condition the model on the vector of TH parameters. In this case, the trained VAE model can only generate synthetic samples randomly, making it difficult to evaluate its performance by comparing with the real experimental data, since it is almost impossible for the random samples to have TH conditions that exactly match the real experimental data.
The model architecture was optimized by performing hyperparameter tuning for 400 iterations. Both the encoder and decoder networks were constructed with six fully connected layers. Training was conducted over around 360 epochs with a batch size of 52. Prior to training, the data underwent shuffling and standardization. The model was trained on 80% of the NRC CHF dataset, while the remaining 20% was reserved for the testing and validation, where the validation set was used for hyperparameter tuning. TensorFlow [31] was used to build the CVAE model. Once the model was trained, the decoder was used to generate CHF values under specific TH conditions in the testing dataset. Subsequently, the generated CHF values were compared with the corresponding experimental CHF values in the testing dataset to evaluate their accuracy.
@@ 4.2 The DNN Predictive Model
The DNN predictive model was also constructed in TensorFlow using eight hidden layers to transform the five input parameters listed in Section 4.1. The network depth, number of neurons per layer, activation functions, and learning rate were optimized using RayTune [32] using 1,000 hyperparameter configurations in a random search. This tuning process was performed using the validation partition of the dataset. The training, validation, and testing datasets used between the CVAE and DNN were identical to ensure an accurate and fair comparison between the two methods. Once the hyperparameters were set, the model was trained for a total of 500 epochs with an exponential learning decay rate of 0.96. The fully trained network was then evaluated against the testing set using the set of metrics provided in Section 5. The DNN model is much more straightforward to train since it only builds a black-box surrogate model of the training dataset without considering the data distribution. Our goal is to use such a fine tuned DNN model to “benchmark” the performance of the CVAE generative model using a comprehensive collection of metrics presented in the following sections.
@@ 4.3 UQ of the ML Models
Quantifying uncertainties in the generated samples by the CVAE model was done by utilizing the inherent variability in the generation process. Recall that the latent space in the CVAE model consists of two vectors that represent the mean value and standard deviation of the encoded distribution, which is usually assumed to be Gaussian. We first obtained 500 samples from the latent space distribution. Each of the latent vector samples is then combined with an input TH condition vector (the conditioning vector $\mathbf{c}$). The combined vector will be fed into the decoder to produce a CHF value. A different CHF value will be generated each time the decoder takes the same TH condition but different latent vector. This process was repeated for all TH conditions in the testing dataset. Afterward, the mean $\mu_{\text{samples}}$ and standard deviation $\sigma_{\text{samples}}$ were computed among the 500 samples for each TH condition. $\sigma_{\text{samples}}$ will be used to represent the uncertainty in the generated samples. It also indicates how sensitive the model predictions are to changes in the latent vector.
@@ 4.4 Domain Generalization
An important factor in ML model evaluation is assessing the model’s ability to generalize to data outside the training domain. This involved assessing how accurately can the model extrapolate to new regions with unseen data. To this end, we employed the concept of a convex hull to evaluate both the CVAE and DNN models. The convex hull is the smallest convex set that includes all points in a given dataset. In order to determine whether the point we are predicting at is inside or outside the training domain, we first construct a convex hull based on the training dataset. Afterward, the points in the testing dataset are classified into two categories; inside the convex hull or outside the convex hull. This is done by examining each testing data point to ascertain whether it falls within or outside the convex hull. If inside, the points are deemed to be within the domain covered by the training data; if outside, then the point is not represented in the space covered by the training data.
In this study, our testing dataset comprises 2,458 data points (10% of the NRC CHF dataset). Of these, 2,307 points fall within the convex hull defined by the training dataset, while 151 points lie outside. Evaluating the two models’ performance on generated/predicted CHF values for the 151 points outside the convex hull allows us to assess its capability for domain generalization.
@@ 5.1 Results of CHF Generation and Prediction using CVAE and DNN
Table 1 includes statistical metrics calculated based on the absolute relative errors and the $R^{2}$ metric for both models. It provides the mean, maximum, and standard deviations of the absolute relative error values, along with the fraction of testing points resulting in an absolute relative error value greater than 10%. The statistical metrics show that both models were able to predict CHF values accurately, with mean absolute relative error values of 7.292% for the CVAE model and 8.455% for the DNN model. The maximum error values were significantly higher than the mean values, however, these are observed at comparatively small CHF values, which can result in larger relative errors despite small absolute differences. Notably, only 6.468% of the testing data points had an absolute relative error greater than 25% for the DNN model, compared to 4.353% for the CVAE model.
@@
$\mu_{\text{error}}$ 8.455 % 7.292 %
$\text{Max}_{\text{error}}$ 94.40 % 113.92 %
$\text{Std}_{\text{error}}$ 10.607 % 9.329 %
$F_{\text{error}}>10\%$ 26.89 % 23.79 %
$\text{R}^{2}$ 0.9845 0.9880
The standard deviation values for both models indicate that the errors are not widely spread from the mean, suggesting that most predictions are close to the actual CHF values. Furthermore, both networks are observed with an $R^{2}$ value above 0.98 indicating a very high degree of correlation between true and generated/predicted values. Overall, these results show that both models perform well in generating and predicting CHF values, with the CVAE having more favorable values across all error metrics, except for $\text{Max}_{error}$ which was observed to be slightly higher.
The last row in Figure 1 presents the pair-wise correlations between the measured CHF values and each of the five parameters in the vector of TH conditions. It is desirable to maintain such correlations in the generated and predicted CHF data. The correlations between the predicted and generated CHF values and the TH parameters in the testing dataset were compared with the real data to determine whether the models can learn and preserve these correlations. Figure 4 compares the correlations of the TH parameters with the true CHF values and the generated CHF values using the CVAE model. Similarly, Figure 5 compares the correlations of the TH parameters with the true CHF values and the predicted CHF values using the DNN model. These figures also illustrate the TH parameter ranges in the testing dataset.
@@
Figures 4 and 5 show that the correlations between the TH parameters and the generated/predicted CHF values align closely with those found in the real data. Both models have consistent behavior in terms of the correlations. While largest relative absolute error is observed at smaller CHF values, the largest absolute deviations from the true values are observed at higher CHF values, which are associated with smaller values of $D$, $L$, and $P$, as well as subcooled coolant conditions, and high $G$ values. It is important to note that these regions contain fewer data points in both training and testing datasets, resulting in the model’s deviations from the true values under these conditions. However, these deviations are not significant and do not heavily influence the error distributions, as these CHF conditions are less common in practical applications and thus in the CHF dataset.
@@ 5.2 Results of UQ
To quantify the model uncertainty, 500 unique samples were computed for each of the testing set’s input entries using the trained CVAE model. From these, the means and standard deviations of these samples were taken for each of these input entries. The ensemble approach described in Section 4.3 was then implemented using $n=20$ models each initialized with a different random seed. Each of these models were identical with the exception of the initialization. As with the CVAE model, the means and standard deviations were computed along the 500 samples for each input vector of the testing set. The standard deviations were then taken to compute the relative standard deviation using Equation 3. This metric standardizes the uncertainty measure by expressing the standard deviation as a percentage of the mean, making it easier to interpret and to compare across different scales of the outputs.
@@
This process resulted in a relative standard deviation value for each of the output means, providing quantification of the model’s uncertainty. By comparing the relative standard deviations from the CVAE to those of the ensemble of DNNs, we can assess the effectiveness of each approach in capturing the uncertainty associated with the model predictions. This is done using the 6 metrics reported in Table 2, which also include the standard deviation of the error across all prediction means, as was done in Section 5.1. In terms of error values, the CVAE achieves a smaller mean relative error of 7.26% when compared to the 7.93% of the CVAE. The standard deviation of this error distribution is also tighter in the CVAE’s case, 9.28% versus 10.42%, but with generally comparable values when considering their small magnitudes. The DNN’s maximum value of the error distribution, 118.50%, is observed to be slightly larger than the CVAE’s of 111.39%. Both of these maximum values are located at relatively small CHF values near the origin.
@@
$\mu_{\text{error}}$ 7.938 % 7.264 %
$\text{Max}_{\text{error}}$ 118.50 % 111.39 %
$\text{Std}_{\text{error}}$ 10.424 % 9.2841%
Mean relative std 4.8240 % 0.6187 %
Max relative std 31.882 % 12.142 %
$F_{\text{error}}>10\%$ 26.078 % 23.759 %
Now considering the distribution of relative standard deviations, the CVAE’s results report significantly smaller mean and maximum relative Std values compared to the DNN’s. The mean relative standard deviation of the DNN (4.82%) is over seven times greater than that of the CVAE (0.6187%), indicating that the CVAE is more confident in its predictions with less variability in its outputs. This is similarly mirrored in their maximum values, where the DNN has a value of nearly 31.88 % with the CVAE reporting 12.14 %. The final metric, the fraction of predictions with relative errors above 10%, has comparable values between the two models. Considering the combination of the metrics in Table 2, the CVAE outputs are shown to have similar performance compared to a DNN ensemble while achieving more favorable values across all error metrics and significantly smaller model uncertainties.